OpenStudy (anonymous):

Power and work \[ W= \int P dt = \int Fv dt= m \int \frac{dv}{dt} vdt=m \int vdv= \frac{1}{2}m(v_1^2-v_0^2) \] Is this true?: -Did I oversimplify somewhere? -Is this unphysical?

OpenStudy (anonymous):

\[ W= \int\limits P dt = \int\limits Fv dt= m \int\limits \frac{dv}{dt} vdt=m \int\limits vdv= \frac{1}{2}m(v_1^2-v_0^2) \]

OpenStudy (anonymous):

it's fine.

OpenStudy (anonymous):

correct

OpenStudy (kainui):

Similarly, \[W=\int\limits_{}^{}Fdx=\int\limits_{}^{}madx=m \int\limits_{}^{} \frac{ dv }{ dt } dx=m \int\limits_{}^{} \frac{ dv }{ dx }\frac{ dx }{ dt } dx = m \int\limits_{}^{} v dv=\frac{ 1 }{ 2 }mv^2\]

OpenStudy (anonymous):

@Kainui That's what I did.

OpenStudy (kainui):

No, you did something slightly different. You used power, I did not.